### v1

# SADDE: Semi-supervised Anomaly Detection with Dependable Explanations

Abstract.
Semi-supervised learning holds a pivotal position in anomaly detection applications, yet identifying anomaly patterns with a limited number of labeled samples poses a significant challenge. Furthermore, the absence of interpretability poses major obstacles to the practical adoption of semi-supervised frameworks. The majority of existing interpretation techniques are tailored for supervised/unsupervised frameworks or non-security domains, falling short in providing dependable interpretations. In this research paper, we introduce SADDE, a general framework designed to accomplish two primary objectives: (1) to render the anomaly detection process interpretable and enhance the credibility of interpretation outcomes, and (2) to assign high-confidence pseudo labels to unlabeled samples, thereby boosting the performance of anomaly detection systems when supervised data is scarce. To achieve the first objective, we devise a cutting-edge interpretation method that utilizes both global and local interpreters to furnish trustworthy explanations. For the second objective, we conceptualize a novel two-stage semi-supervised learning framework tailored for network anomaly detection, ensuring that the model predictions of both stages align with specific constraints. We apply SADDE to two illustrative network anomaly detection tasks and conduct extensive evaluations in comparison with notable prior works. The experimental findings underscore that SADDE is capable of delivering precise detection results alongside dependable interpretations for semi-supervised network anomaly detection systems. The source code for SADDE is accessible at: https://github.com/M-Code-Space/SADDE.
Dependable explanation, semi-supervised learning, anomaly detection
1. Introduction
Anomaly detection has become a fundamental task in various applications, including network intrusion detection (Mirsky et al., 2018), web attack detection (Tang et al., 2020), and advanced persistent threat detection (Bowman et al., 2020). Anomaly detection systems aim to identify unforeseen threats, such as zero-day attacks. To achieve this, semi-supervised learning is increasingly promising, as it requires only limited labeled data, unlike traditional supervised approaches that depend on extensive labeled samples for training.
Despite the great promise of semi-supervised network anomaly detection systems, the lack of interpretability of their predictions poses major barriers to their practical adoption. Firstly, it is challenging to build trust in the decisions made by these systems, which often provide only binary outputs (benign or malicious) without adequate justification or reliable evidence. Second, it is nearly impossible for analysts to manually interpret the decisions of these black-box machine learning models, as the volume of training data and the complexity of the learned models are beyond human understanding (Guidotti et al., 2018b). Third, minimizing false positives and miss-detected samples remains a critical challenge for anomaly detection systems in real-world applications. Without a clear understanding of how the models operate, it is impossible to effectively update or adjust them to reduce them. Consequently, security operators are left unable to decide whether they can trust model decisions based on excessively simple model predictions, are unwilling to use theoretical semi-supervised network anomaly detection systems in their practical applications, and feel overwhelmed by numerous meaningless false positives.
Figure 1. Illustration of the divergent explanation results of different interpreters.
Figure 2. Overview of the proposed SADDE.
Recently, some research has proposed techniques for interpreting machine learning model decisions by identifying key features that significantly influence the final prediction (Montavon et al., 2018; Ribeiro et al., 2016b; Lundberg and Lee, 2017; Guo et al., 2018). Predominant approaches include local/model-agnostic (Ribeiro et al., 2016a, 2018; Guidotti et al., 2018a; Lundberg, 2017; Guo et al., 2018) and global (Fisher et al., 2019; Kauffmann et al., 2024) methods that focus on interpreting individual and clusters predictions for a given black-box classifier. These interpreters either try to approximate the local decision boundary using a linear algorithm to find key features of the given input, or they perturb the current input’s feature values or permute features of a cluster and observe the model’s feedback to pinpoint the most influential ones. Security operators can use these interpreted key features to better understand model decisions (Guidotti et al., 2018b). Existing work like (Fong and Vedaldi, 2017; Shrikumar et al., 2017a; Zhou et al., 2016; Zintgraf et al., 2017; Kauffmann et al., 2024; Mothilal et al., 2020; Shrikumar et al., 2017b) predominantly focuses on interpreting supervised/unsupervised models in non-security domains, while some existing interpreters like (Han et al., 2021; Liu et al., 2017a; Yang et al., 2021; Wei et al., 2023) are proposed for security-related applications. However, these studies cannot be directly used for semi-supervised network anomaly detection. Furthermore, studies indicate that the explanation results provided by existing interpreters lack consensus (Fan et al., 2020), leading to confusion regarding their trustworthiness. To our best knowledge, although a few studies examine the reliability of existing interpreters (Fan et al., 2020; Yeh et al., 2019; Camburu et al., 2019), they mainly focus on developing more effective evaluation metrics rather than improving the reliability of interpreters’ results. Fig. 1 illustrates the process of a classifier analyzing network traffic data and producing decisions, such as identifying whether the traffic is anomalous or benign. Various interpretability methods, including LIME, DeepLIFT, Lemna, and so on, generate different key features (e.g., ”Flow Bytes/s,” ”Idle Max,” ”Fwd IAT Min”), leading to confusion for the security analyst. The inconsistency in these explanations raises uncertainty regarding the reliability of the interpretations.
Our Work. In this paper, we propose SADDE, a general semi-supervised framework for automatically learning from enormous unlabeled data and improving the reliability of interpretation results. The design goal of SADDE is to develop a novel semi-supervised learning framework for network anomaly detection applications that meets the special requirements of security domains (such as reliable, human-understandable, stable, robust, and fast). To this end, we propose two techniques in SADDE referred to as Global-local Knowledge Association Mechanism (KAM) and Two-stage Semi-supervised Learning System (ToS). KAM helps security analysts understand the underlying reason for model predictions while ToS provides an effective pseudo-labeling process to assist the semi-supervised learning. Compared to original black-box machine learning models, security practitioners can better understand system feedback. Besides, with more reliable interpretations, system operators are more willing to adopt theoretical semi-supervised frameworks in their practical applications. We also provide prototype implementations of SADDE over two representative network anomaly detection datasets (ISCXTor2016 (Lashkari et al., 2017) and CIC-DoHBrw-2020 (MontazeriShatoori et al., 2020)). The proposed KAM and ToS are extensively evaluated compared to representative prior approaches. Experimental results demonstrate that SADDE can provide reliable explanations for semi-supervised learning frameworks while satisfying the requirements of security-related applications.
Contributions. This paper has the following contributions:
  • We propose SADDE, a general framework for interpreting and improving semi-supervised network anomaly detection in security-related applications, which includes two key components: KAM provides reliable, human-understandable, stable, robust, and fast interpretations for semi-supervised systems and ToS makes it possible to automatically learn from large unlabeled datasets.
  • We conduct extensive experiments to demonstrate that the proposed SADDE outperforms existing approaches regarding fidelity, stability, robustness, and efficiency.
2. Framework
2.1. Overview
The overview of SADDE is illustrated in Fig. 2 consisting of two key components: Global-local Knowledge Association Mechanism (KAM) and Two-stage Semi-supervised Learning System (ToS). ToS enables accurate semi-supervised learning via a novel two-stage pseudo-labeling process while KAM provides reliable interpretations for the semi-supervised anomaly detection system. SADDE is a general model-independent framework that formulates semi-supervised learning and decision interpretation into a unified end-to-end learning process with security-related constraints. We instantiate SADDE on the aforementioned network anomaly detection security application. With SADDE, malicious network traffic can be automatically detected without requiring large labeled datasets, meanwhile, security practitioners can better understand system behaviors and trust system decisions.
Working Process. The working process of the proposed SADDE is as follows. Before starting, we assume that there is a small labeled dataset that can be utilized to pre-train the local and global machine learning models. Such assumption is reasonable, as semi-supervised learning necessitates the model to continuously assign pseudo-labels to unlabeled samples, and pre-training the model using a small labeled dataset saves data labeling costs while ensuring the effectiveness of the subsequent ”data-labeling” process. Similar assumptions can be found in (Yao et al., 2018; Apruzzese et al., 2022; Li et al., 2024). Given an instance (or a batch of instances), we fed it into the pre-trained local model to extract its feature embedding and predict its label. The extracted feature embedding is then clustered by the global model into several clusters, where samples from the same cluster share similar characteristics. Feature representations from a cluster are treated as global representations while the feature embedding from an instance is defined as a local representation. We group the input instance into one of the clusters by comparing the distance between the embedding of the instance and the clusters. Then, the predictions of the global and local models are interpreted by the global and local interpreters, and the reliability of the interpretation results is further improved by comparing the global and local interpretations. The underlying reason behind it is that global and local interpreters are heterogeneous interpreters. It has a higher probability of improving its trustworthiness by matching global and local interpretations instead of two local interpreters. For instance, in scenarios where individual samples contain noise or have been maliciously altered, the interpretation results at the local level may be wrong, however, due to the fact that the global and local interpreters are heterogeneous, there is a probability that the interpretation results at the global level is correct. In this way, the interpretations’ reliability could be improved. If the interpreted global and local key features match (the similarity score is higher than a given threshold), we consider the key features interpreted by the interpreters to be credible. Conversely, if they do not match, the samples will be discarded. These samples will not be utilized for model re-training to avoid possible model performance deterioration. Finally, labels predicted by the global and local models are aligned to find the most credible pseudo-labels for the next round’s semi-supervised training. More details about SADDE are presented in Algorithm 1.
Input :  Samples $X=\{x_{1},\cdots,x_{N}\}$, local model $M_{1}$ and global model $M_{2}$, local interpreter $L$, and global interpreter $G$
Output : Final predictions and interpreted key features of each sample and each class. Samples with high-confidence pseudo labels. Well-trained $M_{1}$ and $M_{2}$.
% Step 0: Model Pre-training
Train $M_{1}$ and $M_{2}$ on the labeled pre-training set
while True do
       for Each sample (or batch of samples) in $X$ do
             % Step 1: Feature Extraction
             Extract features of $x_{i}$ and predict its local pseudo label $\hat{y}_{1}$ by $M_{1}$
             Group the feature embedding into a cluster by $M_{2}$ using
             $c^{(i)}=\mathop{\arg\min}\limits_{j}||v_{i}-\mu_{j}||^{2}$.
             Assign the label of the $j$-th cluster $c^{(i)}$ to the unlabeled data $x_{i}$, i.e., the global pseudo label $\hat{y}_{2}$.
             % Step 2: Double Verification
             //First Verification: KAM//
             Interpret key features of the cluster by $G$
             Interpret key features of the instance by $L$
             Calculate a similarity score $sim$ between the interpreted global and local key features by
             $sim(int_{x}(m),int_{y}(m^{\prime}),k)=2*\frac{int^{k}_{x}(m)\cap int^{k}_{y}(m^{\prime})}{\left|int^{k}_{x}(m)\right|+\left|int^{k}_{y}(m^{\prime})\right|}.$ if The $sim>T$ then
                   //Second Verification: ToS//
                   Check the predictions of $M_{1}$ and $M_{2}$
                   if $\hat{y}_{1}==\hat{y}_{2}$ then
                        Treat them as samples with reliable pseudo-labels and use them in the next round of model training
                   end if
                  Final output
             end if
       end for
      Retrain $M_{1}$ and $M_{2}$ on the pseudo-labeled data
end while
Algorithm 1 SADDE
2.2. Global-local Knowledge Association Mechanism (KAM)
The global-local Knowledge Association Mechanism (KAM) takes advantage of both the local and global interpreters to provide more reliable interpretations of model predictions. It uses a local (or model-agnostic) interpreter $L$ to interpret the prediction of the local model $M_{1}$. The interpretation results indicate which features influence the model’s prediction the most. Meanwhile, it adopts a global interpreter $G$ to explain the prediction of the global model $M_{2}$, and the interpreted key features represent the most important features of the cluster that the current instance belongs to.
Based on the results obtained by the global interpreter $G$ and local interpreter $L$, a feature similarity calculation method (as defined in Eq.(1)) is used to further verify the reliability of the interpretation results. Inspired by (Fan et al., 2020), the similarity score is defined as follows. Different from (Fan et al., 2020), it measures the similarity between the locally interpreted key feature set of an instance and the globally interpreted key feature set of a cluster that the instance belongs to.
(1) $sim(int_{x}(m),int_{y}(m^{\prime}),k)=2*\frac{int^{k}_{x}(m)\cap int^{k}_{y}(m^{\prime})}{\left|int^{k}_{x}(m)\right|+\left|int^{k}_{y}(m^{\prime})\right|},$
where $int_{x}(m)$ is the interpretation result of the $x$-th sample generated by the interpreter $m$, and $k$ represents the interpreted top-$k$ key features of the $x$-th sample generated by the interpreter $m$. The $int_{y}(m^{\prime})$ has a similar meaning where $m^{\prime}$ is another interpreter. Similarly, $int_{x}^{k}(m)$ represents the number of top-$k$ interpretation results of the $x$-th sample interpreted by interpreter $m$.
The local interpretations of an instance are considered highly reliable when the feature similarity of the local and global interpretations is higher than a given Threshold. In our case study, we utilize the DeepLift (Shrikumar et al., 2017b) as our local interpreter and PFIE (Fisher et al., 2019) as our global interpreter as an example. Other local/model-agnostic or global interpreters can also be used as the local and global interpreters in KAM. Note that different from existing interpreters such as (Han et al., 2021; Liu et al., 2017a; Yang et al., 2021), which only interpret anomalous samples or outliers, SADDE interprets both normal and anomalous samples for that both normal and anomalous samples can potentially be used to update the model in SADDE, thereby influencing the entire learning and decision-making process of anomaly detection. Besides, in security applications, a transparent decision-making process is crucial for system operators of security applications to adopt theoretical algorithms in practice.
2.3. Two-stage Semi-supervised Learning (ToS)
We develop a two-stage semi-supervised learning framework named ToS to effectively assign high-confidence pseudo labels to the unlabeled samples (as illustrated in Fig. 2).
Typically, for semi-supervised learning, pseudo labels are created by a pre-trained machine learning model. In contrast with most existing semi-supervised learning methods that explore pseudo labels by only once predictions, ToS executes this process twice and the confidence of these virtual labels is improved by comparing predictions of both stages. For anomaly detection systems in security domains that only have limited labeled data, ToS repeatedly adds highly confident pseudo-labeled data for re-training the model, which significantly improves the model performance. The key idea of ToS lies in the selection of high-confidence pseudo-labeled samples and thus guarantees the subsequent process of model re-training.
As shown in Fig. 2, ToS contains a local machine learning model $M_{1}$ and a global model $M_{2}$. During the pertaining phase, $M_{1}$ and $M_{2}$ are pre-trained on a smaller labeled set, making them capable of providing virtual labels to unlabeled data. During the re-training phase, firstly, $M_{1}$ extracts features of given instances and makes predictions. Secondly, $M_{2}$ clusters the feature embedding into several clusters according to feature distances (see Eq.(2)).
(2) $c^{(i)}=\mathop{\arg\min}\limits_{j}||v_{i}-\mu_{j}||^{2}$
where $v_{i}$ denotes feature embedding of an unlabeled instance $x_{i}$, and $\mu_{j}$ represents centroids of class $j$ in labeled data. $c^{(i)}$ is the cluster that has the closest distance from $v_{i}$ among all cluster centroids in the labeled data.
Double Verification Mechanism. The goal of the double verify mechanism is to further reduce false positive and miss-detection rates and improve detection accuracy through a two-layer verification process that incorporates both the similar feature verification in KAM and the global and local prediction check in ToS. For the first-layer verification, as aforementioned in Section 2.2, it compares the calculated similarity score (see Eq.(1)), denoted as $sim$, with a predefined threshold $T$. We consider the interpretations to be reliable if $sim>T$. To some extent, it indicates that the sample has not been modified or attacked and could be used for the next round of re-training. For the second-layer verification, we consider the generated pseudo labels to be confident when the local model’s prediction $\hat{y}_{l}$ matches the global model’s prediction $\hat{y}_{g}$. The selected reliable samples and their pseudo-labels are used to retrain both the global and clustering models. The model then waits for new network flow data to arrive for further processing.
In our case study, we use the Multilayer Perceptron Network (MLP) (Rumelhart et al., 1986) and DBSCAN (Ester et al., 1996) as our local and global models respectively as an example for evaluating the effectiveness of the proposed ToS. Other machine learning models can also be chosen accordingly for specific tasks.
3. Evaluation
This section presents the experimental setup and results. The datasets are detailed in subsection 3.1. Next, the experiment setup is introduced in subsection 3.2. Finally, in subsection 3.3, we compare the proposed KAM and ToS with five and eight state-of-the-art on two representative network anomaly detection datasets and analyze their results.
3.1. Dataset
Two widely used network anomaly detection datasets, ISCXTor2016 (Lashkari et al., 2017) and CIC-DoHBrw-2020 (MontazeriShatoori et al., 2020), are employed to evaluate the performance of SADDE. ISCXTor2016 contains Tor and non-Tor network flows in various applications and each instance has 28 features. It consists of 68,191 samples, of which 56,534 are non-Tor and 11,657 are Tor samples. CIC-DoHBrw-2020 contains labeled traffic generated by DNS over HTTPS (DoH) protocols. It is a dataset for detecting browser-based DoH activities and studying secure communication anomalies. Each instance has 33 features. There are 933,189 instances, including 216,649 DoH and 716,540 non-DoH samples.
For the ISCXTor2016 dataset, the first 1% of the data is used as the pre-training set (i.e., small labeled set), the last 20% as the testing set, and the next 20% as the validation set, while the remaining data is allocated for re-training. The CSIC-DoHBrw-2020 dataset follows a similar partitioning structure, except that the first 0.5% of the data is used as the pre-training set.
3.2. Setup
System Settings. The experiments are conducted using Python (v3.8), along with the PyTorch framework (v1.11.0), CUDA (v11.3), and Scikit-learn (v1.3.2) to develop and evaluate SADDE. A three-layer FCN is employed as our $M_{1}$, and each layer contains 256 neurons.
During training, the Adam optimizer is used with a learning rate of $31e-$ 3. A dropout rate of 0.3 is implemented to avoid over-fitting, and cross-entropy loss was selected as the objective function for classification tasks. We set the batch size to 64. DBSCAN was employed with an epsilon value of 0.5 and a minimum sample size of 1 for clustering, ensuring that all samples could potentially be included in a cluster. The output of FCN’s third fully connected layer is utilized as the input of DBSCAN. Additionally, early stopping was implemented with a patience of 5 epochs during both the pre-training and re-training phases so that the training process could be halted if there is no improvement in the validation loss over 5 epochs. In PFIE, each feature is permuted by 100 times. For DeepLift, the eps parameter is configured to $101e-$ 10. The number of key features is specified as 10, and the threshold for the similarity score is established at 0.6.
All simulations are executed on a Linux server powered by an AMD EPYC 7282 16-Core Processor (2.80 GHz) and 377 GB of RAM. To ensure consistency, all algorithms are run within the same environment, and for fairness, the average results from 10 runs are recorded for comparison.
Benchmarks. We employ a comprehensive set of benchmarks in this work to validate the performance of SADDE. Firstly, we conduct a comparative analysis of the proposed KAM against eight state-of-the-art interpreters, including Lime (Ribeiro et al., 2016b), SHAP (Lundberg and Lee, 2017), DeepLift (Shrikumar et al., 2017b), COIN (Liu et al., 2017b), DiCE (Mothilal et al., 2020), xNIDS (Wei et al., 2023), CADE (Yang et al., 2021), Anchors (Ribeiro et al., 2018).
  • Lime  [^1] https://github.com/marcotcr/lime: It explains the predictions of any classifier by learning an interpretable model in the local vicinity of the prediction.
  • SHAP  [^2] https://github.com/shap/shap: It assigns each feature an important value using principles from cooperative game theory.
  • DeepLift  [^3] https://github.com/pytorch/captum: It is an interpreter that decomposes a neural network’s output prediction for a specific input by backpropagating the contributions of all neurons in the network to each feature of the input.
  • COIN  [^4] https://github.com/xuhongzuo/outlier-interpretation: It transforms outlier detection into local classification tasks, identifying key features and assigning outlier scores to explain the differences between outliers and their surrounding normal instances.
  • DiCE  [^5] https://github.com/microsoft/DiCE: It generates diverse counterfactual explanations by optimizing for both proximity to the original input and diversity among examples.
  • xNIDS  [^6] https://github.com/CactiLab/code-xNIDS: It approximates and samples around historical inputs, and captures feature dependencies of structured data to provide high-fidelity explanations.
  • CADE  [^7] https://github.com/whyisyoung/CADE: It identifies features that cause the largest changes in distance between a drifting sample and its nearest class within a low-dimensional latent space learned through contrastive learning.
  • Anchors  [^8] https://github.com/SeldonIO/alibi: It provides rules that sufficiently ”anchor” the prediction locally, ensuring that changes to the rest of the feature values do not affect the prediction.
Then, we compared the introduced ToS with five state-of-the-art semi-supervised methods, including InstantT (Li et al., 2024), ACR (Wei and Gan, 2023), Clustering* (Fini et al., 2023), LVM (Min et al., 2024), and M3S (Sun et al., 2020a).
  • InstantT  [^9] https://github.com/tmllab/2023_NeurIPS_InstanT: It assigns pseudo labels to unlabeled data by using a threshold-based approach that dynamically adjusts instance-dependent thresholds in response to label noise and prediction confidence.
  • ACR  [^10] https://github.com/Gank0078/ACR: It dynamically refines pseudo-labels across various distributions by estimating the true class distribution of unlabeled data using a unified equation.
  • Clustering*: It utilizes a multi-task framework that combines a supervised objective using ground-truth labels with a semi-supervised objective based on clustering assignments, and optimized through a single cross-entropy loss.
  • LVM: It incorporates the dimension of local variance into pseudo-label selection.
  • M3S: It is a multi-stage training framework that combines the DeepCluster technique with an alignment mechanism in the embedding space to enhance the training process.
Evaluation Metrics. We use Accuracy (Acc.), Precision (Pre.), Recall (Rec.), F1-score (F1.), False Alarm Rate (FAR), Missed Detection Rate (MDR), and standardized partial AUC (SPAUC) (McClish, 1989) to measure different semi-supervised models’ performance. Particularly, SPAUC measures the performance of a model within a specific region of the ROC curve, focusing on a segment that is most relevant under conditions of class imbalance and differential costs of misclassification. FAR reflects the amount of normal network traffic that is wrongly classified as abnormal while MDR indicates how much anomaly network traffic that are miss-detected by the models. The higher the SPAUC, the better the performance is. The lower the FAR and MDR, the better the performance is. Here we use $\textit{SPAUC}_{\textit{FPR}\leq 0.05}$ for all experiments.
Additionally, we evaluate the performance of the proposed KAM interpreter in terms of fidelity, stability, robustness, efficiency, and AOPC, and compare it with the state-of-the-art interpreters.
Figure 3. Fidelity, stability, and robustness evaluation of interpreters on ISCXTor2016.
Figure 4. Fidelity, stability, and robustness evaluation of interpreters on CIC-DoHBrw-2020.
  • Fidelity. To evaluate the fidelity of an interpreter, we define an indicator similar to that used in (Han et al., 2021), called the Label Flipping Rate (LFR), which is the ratio of samples that change to a different class after being modified based on the interpretation results. Specifically, in the context of a classification task, the interpreter identifies several key features. The values of these key features are replaced with the average values corresponding to the opposite class, and a new prediction is generated to determine whether the label has been changed. The LFR is mathematically expressed as:
    (3) $LFR=\frac{\sum_{i=1}^{n}I(f^{\prime}(x_{i})\neq f(x_{i}))}{n}$
    where $n$ denotes the total number of samples, $f(x_{i})$ represents the original predicted label for a sample $x_{i}$, $f^{\prime}(x_{i})$ is the predicted label after modification, and $I(\cdot)$ is an indicator function that equals 1 if the label has flipped and 0 otherwise. This metric captures the percentage of samples for which the predicted label changes after altering the key features.
  • Stability. The stability of interpreters refers to the consistency of interpretation results for the same samples across multiple runs. Given two interpretation outcomes $int_{x}(m)_{1}$ and $int_{x}(m)_{2}$, their similarity can be calculated using Eq.(1). This process is repeated to assess the similarity between the interpretation results from two runs under identical settings for each sample, and the overall stability is calculated as follows:
    (4) $Stability=\frac{1}{n}\sum_{i=1}^{n}\text{sim}(int_{x}(m)_{1}^{i},int_{x}(m)_{2}^{i})$
    In this equation, $n$ is the total number of samples, and $\text{sim}(int_{x}(m)_{1}^{i},int_{x}(m)_{2}^{i})$ represents the similarity between the two interpretation outcomes for the same sample $i$. The average similarity across all samples is calculated to quantify the stability of the interpreter.
  • Robustness. Robustness refers to the similarity of interpretation results for the same sample before and after adding noise, sampled from a Gaussian distribution $N(0,\sigma^{2})$. Specifically, for each sample $i$, the robustness is calculated as the average similarity between the interpretation result without noise, $int_{x}(m)_{i}$, and the interpretation result after adding noise, $int_{x}(m+\epsilon)_{i}$, where $\epsilon\sim N(0,\sigma^{2})$. The overall robustness is computed using the equation:
    (5) $Robustness=\frac{1}{n}\sum_{i=1}^{n}\text{sim}(int_{x}(m)_{i},int_{x}(m+\epsilon)_{i},k)$
    Here, $k$ represents the number of key features. In this work, the noise is set to $\sigma=0.1$.
  • Efficiency. We assess efficiency by recording the runtime required to interpret 2000 samples for each interpreter.
3.3. Results and Analysis
Verification of Model Pre-training. Table 1 and Table 2 present the performance of ToS when trained on the pre-training set (denoted as ①) and when trained on both the pre-training and re-training sets (denoted as ②). The results indicate that the re-training approach (②) achieves significantly better performance than pre-training alone (①), with improvements of approximately 20%, 10%, and 14% in Precision, FAR, and SPAUC, on the ISCXTor2016 dataset. Similarly, on the CIC-DoHBrw-2020 dataset, the re-training process (②) enhances the model’s performance by about 14%, 26%, and 4% in terms of F1-score, MDR, and SPAUC, compared to training only on the smaller pre-training set (①). These findings demonstrate the effectiveness of the semi-supervised learning process in improving the performance of the proposed SADDE.
Table 1. Performance (%) of ToS on ISCXTor2016: comparisons between model performance (%) with only the pre-training set (①) and with both pre-training and re-training set (②). $\triangledown$ denotes the performance difference relative to the former. The best metric performance is bolded.
Setting Acc. Pre. Rec. F1. FAR MDR SPAUC
80.88 65.68 67.04 66.29 12.47 53.44 53.50
89.43 85.36 71.97 76.27 2.18 53.88 67.42
$\triangledown$ +8.55 +19.68 +4.93 +9.98 -10.29 -0.44 +13.92
Table 2. Performance (%) of ToS on CIC-DoHBrw-2020: comparisons between model performance (%) with only the pre-training set (①) and with both pre-training and re-training set (②). $\triangledown$ denotes the performance difference relative to the former. The best metric performance is bolded.
Setting Acc. Pre. Rec. F1. FAR MDR SPAUC
74.67 65.68 67.36 66.35 18.99 46.27 52.34
84.44 78.19 82.86 79.95 14.18 20.07 55.94
$\triangledown$ +9.77 +12.51 +15.50 +13.60 -4.81 -26.20 +3.60
Figure 5. Efficiency evaluation of interpreters.
Figure 6. Performance of ToS compared with the state-of-the-art semi-supervised models: F1-score vs. SPAUC.
Figure 7. Performance of ToS compared with the state-of-the-art semi-supervised models: FAR vs. MDR.
Comparison with the State-of-the-art Interpreters. Figures 3 and 4 illustrate the evaluation of various interpreters across the ISCXTor2016 and CIC-DoHBrw-2020 datasets, highlighting performance differences across three critical dimensions: fidelity, stability, and robustness. Fig. 5, in turn, presents the results of the efficiency evaluation. KAM demonstrates a superior balance across all these dimensions compared to other methods. Notably, while COIN outperforms KAM in terms of fidelity and robustness on the ISCXTor2016 dataset, its efficiency is significantly lower, being nearly 100 times slower than KAM.
We observe that in the ISCXTor2016 dataset, the fidelity of all interpreters is relatively low. This is because anomaly samples, such as Tor traffic, are more susceptible to label flipping, as modifications to their key features are more likely to alter model predictions. However, the dataset’s significant class imbalance—reflected in a normal-to-anomalous ratio of approximately 4.85:1—leads to a lower overall LFR. In contrast, the CIC-DoHBrw-2020 dataset has a normal-to-anomalous ratio of about 3.31:1, offering a relatively more balanced scenario for evaluating fidelity.
According to Fig. 5, KAM demonstrates a balanced performance between interpretability, reliability, and efficiency. While DeepLift and xNIDS are more efficient compared to our method, they do not perform as well on other key metrics. Moreover, we observed that Anchors is less efficient on the CIC-DoHBrw-2020 dataset compared to the ISCXTor2016 dataset. This is due to the combinatorial optimization involved in the search for an anchor, which causes a significant drop in efficiency as the number of features increases. With 28 features in the ISCXTor2016 dataset and 33 in the CIC-DoHBrw-2020 dataset, this increase in feature count directly contributes to the observed difference in performance.
Comparison with the State-of-the-art Semi-supervised Models. As shown in Table 3 and Table 4, ToS outperforms other methods. ToS achieves SPAUC as high as 67.42% and 55.94% on ISCXTor2016 and CIC-DoHBrw-2020 datasets with about 1.33% and 1.53% improvement, respectively. Notably, LVM performs similarly to ToS on the ISCXTor2016 dataset, with only a 0.82% difference in f1-score and 1.33% difference in SPAUC. However, its performance on the CIC-DoHBrw-2020 dataset is significantly lower than ToS, with a 19.47% gap in f1-score and 4.48% in SPAUC, indicating that ToS exhibits better stability and is more adaptable to different scenarios.
Furthermore, as illustrated in Fig. 7, ToS strikes the optimal balance between FAR and MDR. Although the Clustering* method exhibits an exceptionally low FAR on the ISCXTor2016 dataset, its MDR is notably high, nearing 100%. Similarly, while instantT demonstrates superior FAR performance compared to ToS on the CIC-DoHBrw-2020 dataset, both suffer from significantly elevated MDR values. As shown in Fig. 6, ToS also excels in balancing F1-score and SPAUC, outperforming other models on both two datasets.
Table 3. Performance (%) of ToS compared with the state-of-the-art semi-supervised models on ISCXTor2016.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
instantT 84.30 70.82 69.14 69.91 8.42 53.30 55.83
ACR 70.18 56.84 60.23 56.96 25.04 54.51 51.05
Clustering* 83.05 49.67 49.74 46.32 1.07 98.97 49.98
LVM 88.97 83.62 71.45 75.45 2.63 54.47 66.09
M3S 83.56 69.18 68.50 68.83 9.51 53.48 54.99
ToS (ours) 89.43 85.36 71.97 76.27 2.18 53.88 67.42
Table 4. Comparison with the state-of-the-art semi-supervised models’ performance (%) on CIC-DoHBrw-2020.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
instantT 78.07 68.68 58.40 59.20 4.89 78.32 54.41
ACR 80.61 75.62 83.84 77.15 22.19 10.12 53.91
Clustering* 76.51 72.03 79.67 72.84 26.22 14.43 52.90
LVM 71.81 60.47 60.49 60.48 18.39 60.41 51.46
M3S 79.56 72.94 78.12 74.50 19.19 24.57 53.76
ToS (ours) 84.44 78.19 82.86 79.95 14.18 20.07 55.94
4. Discussion
Although in this work we only evaluated SADDE on network anomaly detection tasks as an example, the proposed KAM and ToS can be utilized in a wide range of semi-supervised anomaly detection applications. Besides, the robustness evaluation of AnoamlyIAP is mainly considered against data poisoning/feature perturbation attacks (See Section 2.2). There are other attacks that may fail SADDE like model poisoning attacks because SADDE is highly dependent on the local and global models’ predictions. Future work can investigate the SADDE’s robustness against more attacks. Moreover, hyper-parameters of SADDE are configured empirically. Future research can design more advanced approaches to configure hyperparameters. In this work, different data distribution imbalance ratio is not experimented due to the page limitation. Research on the influence of different imbalance ratios or the performance of SADDE under extreme data distributions can be conducted in the future. Furthermore, more anomaly detection scenarios with different model combinations can also be tested in the future to explore the generalization performance of SADDE. Finally, it is worthy of exploring how to use SADDE on scenarios with only a few labeled samples, for example, by developing better multi-stage semi-supervised learning approaches or new matching strategies.
5. Related Work
In this section, we briefly discuss existing literature from two aspects.
Interpreters. Numerous studies have been conducted to interpret machine learning models and achieved promising results, which can be roughly categorized into two groups, i.e., local/model-agnostic (Ribeiro et al., 2016a, 2018; Guidotti et al., 2018a; Lundberg, 2017; Guo et al., 2018; Fong et al., 2019) and global interpreters (Fisher et al., 2019; Kauffmann et al., 2024). The former interprets individual predictions for a given machine learning model while the latter attempts to explain model predictions of a cluster. Local/model-agnostic interpreters are mostly perturbation-based, Gradient-based, and others. Perturbation-based approaches (Fong et al., 2019; Fong and Vedaldi, 2017; Yang et al., 2021) perturb data features and check the variation of model predictions to identify the most important ones. Gradient-based approaches (Shrikumar et al., 2017b; Selvaraju et al., 2017) back-propagate gradients through the model to evaluate each feature’s sensitivity. Others treat the target machine learning model as a black-box (Ribeiro et al., 2016b, 2018). Techniques like LIME (Ribeiro et al., 2016b), LIMNA (Guo et al., 2018), and SHAP (Lundberg and Lee, 2017) attempt to utilize a linear model to approximate the decision boundary of the input, and utilize it to find the most influential features. Global interpreters like (Fisher et al., 2019) explain model prediction by randomly permuting specific feature values of a cluster and observer model predictions. The authors of (Kauffmann et al., 2024) developed an interpreter for explaining clustering algorithms by formulating clustering decisions as being functionally equivalent neural networks.
Few studies aim to interpret unsupervised models such as COIN (Liu et al., 2017b) and CADE (Yang et al., 2021) which are used as baselines in our experiments. In addition to them, most of other works directly use the interpreters proposed for supervised learning to unsupervised learning applications.
The work of (Fan et al., 2020) demonstrated that the explanation results provided by existing interpreters lack consensus, resulting in confusion regarding their trustworthiness. There are a few studies that examine the reliability of existing interpreters (Yeh et al., 2019; Camburu et al., 2019; Fan et al., 2020), for example, (Yeh et al., 2019) introduced two metrics to evaluate interpreters, i.e., sensitivity and infidelity. An off-the-shelf evaluation test is proposed for post-hoc interpreters in view of feature selection in (Camburu et al., 2019). (Fan et al., 2020) introduced three evaluation metrics to measure interpreters’ stability, effectiveness, and robustness. The work of (Warnecke et al., 2020) developed criteria for evaluating and comparing different interpreters, including both general properties like accuracy and security-related ones, such as robustness, efficiency, and completeness. Unfortunately, to our best knowledge, there is no existing work exploring how to improve the reliability of interpretation results.
Semi-supervised Learning. A recent surge of interest has focused on anomaly detection with semi-supervised learning, which attempts to assign high-confidence pseudo-labels to unlabeled data. There is a wide range of research on semi-supervised learning, which generally falls into two categories, i.e., single-stage-based and multi-stage-based approaches. Some single-stage-based semi-supervised learning techniques improve the confidence of pseudo labels based on thresholds (Li et al., 2024; Min et al., 2024), and they usually set high thresholds to unlabeled samples to prevent the occurrence of incorrect pseudo labels, or contrastive learning (Fini et al., 2023; Wei and Gan, 2023) that combines combine pseudo labeling and consistency regularization to encourages similar predictions between two different views of an instance to improve the robustness of the model. Unlike single-stage-based semi-supervised learning techniques, multi-stage-based semi-supervised learning approaches, such as (Sun et al., 2020a, b), explore the most confident pseudo labels by using multiple models and improve their confidence by aligning the predictions from each stage.
6. Conclusion
In this paper, we propose SADDE, a general framework aiming to interpret and improve semi-supervised anomaly detection performance. It incorporates two key techniques KAM and ToS: (1) KAM provides reliable, human-understandable, stable, robust, and fast interpretations for semi-supervised systems; (2) ToS makes it possible to automatically learn from large unlabeled datasets. By applying and evaluating Adaptive NAD over two classical network anomaly detection datasets, we demonstrate that SADDE can provide fast and accurate anomaly detection results as well as reliable, human-understandable, stable, robust, and fast interpretations.